OpenStudy (anonymous):

Combining like terms. 7 + 1/2y + 1/3y - 8

OpenStudy (aroub):

1/2y+1/3y-1

OpenStudy (anonymous):

1/2+1/3y =3/6 +2/6= 5/6y - 1 is the answer

OpenStudy (anonymous):

are you sure?(:

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

because 1/2y and 1/3y are like terms so u have to add them

OpenStudy (aroub):

^ yeah nicole is right :) I just didnt simplify it more

OpenStudy (anonymous):

@aroub if u could please help with some of my unanswered questions please no one wants to answer them

OpenStudy (anonymous):

wait never mind i don't have any new ones but i'll put more please help

OpenStudy (anonymous):

Okay thank you guys! :)

OpenStudy (aroub):

Yea sure! I'll try ;)

OpenStudy (aroub):

You're welcome!

OpenStudy (anonymous):

Can you answer this one?(: 1/5 - 3/5 + 3/4m

OpenStudy (anonymous):

3/4m - 4/5

OpenStudy (anonymous):

4/5 or 3/5?

OpenStudy (anonymous):

whoops its actually 2/5

OpenStudy (anonymous):

-3/7x - 2/7x